# On thermal transpiration and thermomolecular pressure difference

### v2

@@ 4.1. The existence of solution to (80)
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq C_{\zeta}\max\{1,\kappa^{-1}\}(\mathcal{A}+\mathcal{B})$
@@
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ (132)
$\leq C_{\zeta}\max\{1,\kappa^{-1}\}\Big[(\mathcal{A}+\mathcal{B})+C\mathcal{B}(\mathcal{A}+\mathcal{B})+4CC_{\zeta}\max\{1,\kappa^{-3}\}(\mathcal{A}+\mathcal{B})^{2}\Big]$ (133)
$\leq C_{\zeta}\max\{1,\kappa^{-1}\}\Big[1+C\mathcal{B}+4CC_{\zeta}\max\{1,\kappa^{-2}\}(\mathcal{A}+\mathcal{B})\Big](\mathcal{A}+\mathcal{B}).$
@@
$\delta<\min\left\{\frac{1}{4C},\frac{1}{16CC_{\zeta}\max\{1,\kappa^{-2}\}}\right\}.$
@@
$\lVert f\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq C\max\{1,\kappa^{-1}\}(\mathcal{A}+\mathcal{B}).$ (134)
@@
Note that the $f$ in 138 has the form
@@
Since $\Omega$ is an open set, there exist $\rho>0$ and $a_{1}=(a_{1,1},a_{1,2},0),a_{2}=(a_{2,1},a_{2,2},L)\in\mathbb{R}^{3}$ such that
@@
$\Pi=B(a_{1},\rho)\cap\{x_{3}=0\}\cup B(a_{2},\rho)\cap\{x_{3}=L\}.$ (156)
We show that $\Omega\cap\{0\leq x_{3}\leq L\}$ is completely unsealed with respect to $\Pi$. We can then deduce the lemma from this conclusion. For any fixed $x\in\Xi\coloneq\partial(\Omega\cap\{0\leq x_{3}\leq L\})\setminus(\{x_{3}=0\}\cup\{x_{3}=L\})$, without loss of generality, we may assume that $a_{1}=(0,0,0)$ and $x=(x_{1},0,x_{3})$. Then we classify $v\in\mathbb{R}^{3}$ as follows.
@@
Let $v=(\rho\sin\theta\cos\phi,\rho\sin\theta\sin\phi,\rho\cos\theta)$. Then the sets $A_{1}$ and $A_{2}$ can be written more explicitly as follow.
$A_{1}=$ $\left\{(\rho,\phi,\theta)\in\mathbb{R}_{+}\times[0,2\pi]\times[0,\frac{\pi}{2}]\,\Big|0\leq\alpha_{1}^{2}+\gamma_{1},\quad\frac{\alpha_{1}-\sqrt{\alpha_{1}^{2}+\gamma_{1}}}{x_{3}}\leq\tan\theta\leq\frac{\alpha_{1}+\sqrt{\alpha_{1}^{2}+\gamma_{1}}}{x_{3}}\right\}$ (157)
$A_{2}=$ $\left\{(\rho,\phi,\theta)\in\mathbb{R}_{+}\times[0,2\pi]\times[\frac{\pi}{2},\pi]\,\Big|0\leq\alpha_{2}^{2}+\gamma,\quad\frac{-\alpha_{2}-\sqrt{\alpha_{2}^{2}+\gamma_{2}}}{L-x_{3}}\leq\tan\theta\leq\frac{-\alpha_{2}+\sqrt{\alpha_{2}^{2}+\gamma_{2}}}{L-x_{3}}\right\}$ (158)
@@
$\begin{aligned} &\alpha_{1}=x_{1}\cos\phi,\,&&\gamma_{1}=\rho^{2}-x_{1}^{2}\\ &\alpha_{2}=[x_{1}-a_{2,1}]\cos\phi-a_{2,2}\sin\phi,\,&&\gamma_{2}=\rho^{2}-[(x_{1}-a_{2,1})^{2}+a_{2,2}^{2}].\end{aligned}$
@@
$W_{1}(x)\coloneq\int_{u\cdot n(x)>0}1_{\partial(\Omega\cap\{0\leq x_{3}\leq L\})\setminus A_{1}}(q(x,u))m_{T(q(x,u))}u\cdot n(x)du=1-\int_{u\cdot n(x)>0}1_{A_{1}}(u)m_{T(q(x,u))}u\cdot n(x)du$
@@
$W_{2}(x)\coloneq\int_{u\cdot n(x)>0}1_{\partial(\Omega\cap\{0\leq x_{3}\leq L\})\setminus A_{2}}(q(x,u))m_{T(q(x,u))}u\cdot n(x)du=1-\int_{u\cdot n(x)>0}1_{A_{2}}(u)m_{T(q(x,u))}u\cdot n(x)du.$
Since $\Omega\cap\{0\leq x_{3}\leq L\}$ is a convex set and $\partial\Omega\cap\{0\leq x_{3}\leq L\}$ is a $C^{1}$ boundary, $W_{1}$ is continuous on $\partial\Omega\cap[\frac{L}{2},L]$ and $W_{2}$ is continuous on $\partial\Omega\cap[0,\frac{L}{2}]$. Moreover, $W_{1}$ and $W_{2}$ are continuous on $\overline{\Omega\cap\{x_{3}=L\}\setminus\Pi}$ and $\overline{\Omega\cap\{x_{3}=0\}\setminus\Pi}$ respectively. Hence, we have
@@
$\leq\max\left\{\max_{x\in\partial\Omega\cap[\frac{L}{2},L]}W_{1}(x),\max_{x\in\overline{\Omega\cap\{x_{3}=L\}\setminus\Pi}}W_{1}(x),\max_{x\in\partial\Omega\cap[0,\frac{L}{2}]}W_{2}(x),\max_{x\in\overline{\Omega\cap\{x_{3}=0\}\setminus\Pi}}W_{2}(x)\right\}$
@@ Acknowledgments
Prof. Tai-Ping Liu kindly discussed part of this problem with us. The authors would like to thank him for his interest and insightful suggestions. Also, the author would like to thank researcher Jhe-Kuan Su for giving fruitful comments. The research of the authors is supported in part by NSTC Grant 112-2115-M-002-009-MY3.

### v3

@@ 4.1. The existence of solution to (80)
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq C_{\zeta}(\mathcal{A}+\mathcal{B})$
@@
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ $\leq C_{\zeta}\Big[(\mathcal{A}+\mathcal{B})+C\mathcal{B}(\mathcal{A}+\mathcal{B})+4CC_{\zeta}\max\{1,\kappa^{-1}\}(\mathcal{A}+\mathcal{B})^{2}\Big]$ (132)
$\leq C_{\zeta}\Big[1+C\mathcal{B}+4CC_{\zeta}\max\{1,\kappa^{-1}\}(\mathcal{A}+\mathcal{B})\Big](\mathcal{A}+\mathcal{B}).$
@@
$\delta<\min\left\{\frac{1}{4C},\frac{1}{16CC_{\zeta}\max\{1,\kappa^{-1}\}}\right\}.$
@@
$\lVert f\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq C(\mathcal{A}+\mathcal{B}).$ (133)
@@
Note that the $f$ in 137 has the form
@@
Since $\Omega$ is an open set, there exist $\rho>0$ and $a_{1}\in\{x_{3}=0\},a_{2}\in\{x_{3}=L\}$ such that
@@
$\Pi\coloneq B(a_{1},\rho)\cap\{x_{3}=0\}\cup B(a_{2},\rho)\cap\{x_{3}=L\}.$ (155)
We show that $\Omega\cap\{0\leq x_{3}\leq L\}$ is completely unsealed with respect to $\Pi$. We can then deduce the lemma from this conclusion. For any fixed $x\in\Xi\coloneq\partial(\Omega\cap\{0<x_{3}<L\})$, without loss of generality, we may assume that $a_{1}=(0,0,0)$, $x=(x_{1},x_{2},x_{3})$, and $a_{2}=(a_{2,1},a_{2,2},L)$. Then we classify $v\in\mathbb{R}^{3}$ as follows.
@@
Let $v=(\hat{\rho}\sin\theta\cos\phi,\hat{\rho}\sin\theta\sin\phi,\hat{\rho}\cos\theta)$. Then the sets $A_{1}$ and $A_{2}$ can be written more explicitly as follow.
$A_{1}=$ $\left\{(\hat{\rho},\phi,\theta)\in\mathbb{R}_{+}\times[0,2\pi]\times[0,\frac{\pi}{2}]\,\Big|0\leq\alpha_{1}^{2}+\gamma_{1},\quad\frac{\alpha_{1}-\sqrt{\alpha_{1}^{2}+\gamma_{1}}}{x_{3}}<\tan\theta<\frac{\alpha_{1}+\sqrt{\alpha_{1}^{2}+\gamma_{1}}}{x_{3}}\right\}$ (156)
$A_{2}=$ $\left\{(\hat{\rho},\phi,\theta)\in\mathbb{R}_{+}\times[0,2\pi]\times[\frac{\pi}{2},\pi]\,\Big|0\leq\alpha_{2}^{2}+\gamma,\quad\frac{-\alpha_{2}-\sqrt{\alpha_{2}^{2}+\gamma_{2}}}{L-x_{3}}<\tan\theta<\frac{-\alpha_{2}+\sqrt{\alpha_{2}^{2}+\gamma_{2}}}{L-x_{3}}\right\}$ (157)
@@
$\begin{aligned} &\alpha_{1}=x_{1}\cos\phi+x_{2}\sin\phi,\,&&\gamma_{1}=\rho^{2}-\left(x_{1}^{2}+x_{2}^{2}\right)^{2}\\ &\alpha_{2}=[x_{1}-a_{2,1}]\cos\phi+[x_{2}-a_{2,2}]\sin\phi,\,&&\gamma_{2}=\rho^{2}-[(x_{1}-a_{2,1})^{2}+(x_{2}-a_{2,2})^{2}].\end{aligned}$
@@
$W_{1}(x)\coloneq=1-\int_{u\cdot n(x)>0}1_{A_{1}}(u)m_{T(q(x,u))}u\cdot n(x)du$
@@
$W_{2}(x)\coloneq 1-\int_{u\cdot n(x)>0}1_{A_{2}}(u)m_{T(q(x,u))}u\cdot n(x)du.$
Since $\Omega\cap\{0\leq x_{3}\leq L\}$ is a convex set and $\partial\Omega\cap\{0\leq x_{3}\leq L\}$ is a $C^{1}$ boundary, $W_{1}$ is continuous on $\partial\Omega\cap\{\frac{L}{2}\leq x_{3}\leq L\}$ and $W_{2}$ is continuous on $\partial\Omega\cap\{0\leq x_{3}\leq\frac{L}{2}\}$. Moreover, $W_{1}$ and $W_{2}$ are continuous on $\overline{\Omega\cap\{x_{3}=L\}\setminus\Pi}$ and $\overline{\Omega\cap\{x_{3}=0\}\setminus\Pi}$ respectively. Hence, we have
@@
$\leq\max\left\{\max_{x\in\partial\Omega\cap\{\frac{L}{2}\leq x_{3}\leq L\}}W_{1}(x),\max_{x\in\overline{\Omega\cap\{x_{3}=L\}\setminus\Pi}}W_{1}(x),\max_{x\in\partial\Omega\cap\{\frac{L}{2}\leq x_{3}\leq L\}}W_{2}(x),\max_{x\in\overline{\Omega\cap\{x_{3}=0\}\setminus\Pi}}W_{2}(x)\right\}$
@@ Acknowledgments
The authors would like to thank professor Kazuo Aoki and professor Tai-Ping Liu for their interest and insightful suggestions. Also, the authors would like to thank professor Chun-Hsiung Hsia and researcher Jhe-Kuan Su for giving fruitful comments. The research of the first author is supported in part by NSTC Grant 112-2115-M-002-009-MY3.